OpenStudy (anonymous):

What is the solution set for |–x + 2| = 5?

OpenStudy (anonymous):

is the "|"sign equal the parenthesis?

OpenStudy (jackellyn):

-x+2=5 -x+2=-5 -x = 3 -x= -7 x=-3 x= 7

OpenStudy (anonymous):

x= -3?

OpenStudy (anonymous):

its 2 things _ and _

OpenStudy (anonymous):

7 and -3 or -7 and -3 are 2 choices though

OpenStudy (anonymous):

7